# Mitigating Heterogeneity in Federated Multimodal Learning with Biomedical Vision-Language Pre-training

### v1

@@ 4.1 Guidance-based Local Client Training
, where $N_{c}$ refers to the number of samples in client the $c$. We combine the guidance loss with the optimization goal $\mathcal{L}_{task}$ of the initial biomedical vision-language pre-training method, which includes e.g. image-text contrastive loss in ConVIRT (Zhang et al., 2022b). The overall loss function for local client training can be written as:
$\footnotesize\mathcal{L}_{local}=\mathcal{L}_{guide}+\alpha\mathcal{L}_{task}$ (2)
@@
where $\alpha$ is the hyper-parameter that balances between guidance regularization and training task in local client updating.
@@
4.2 Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
In order to guide the local client training, we need to provide a teacher alignment module as mentioned in the last section, which can capture globally unbiased and generalizable cross-domain alignment. How can we obtain such a teacher guidance model? We propose to learn the unbiased cross-domain alignment by considering the robustness across various distributions of local clients. Detailed problem formulation and description could be seen in Appendix B. A teacher alignment module should avoid distorting the aggregated feature encoder while learning the unbiased cross-modal alignments. Additionally, we introduce a much harder task to force the module to be more robust to biased alignment, by considering the influence of the biased alignment module on the feature encoder. Furthermore, to learn a teacher model that performs well on varies data distributions, we minimize distortion degree under the most biased local alignment with the “worst-case” local data distribution. Learning robust cross-modal alignment through distributionally robust optimization. Inspired by the distributionally robust optimization (DRO) (Cotter et al., 2019), we model the de-centralized pre-training framework as a two-stage DRO learning process. In the first stage, the local models are trained with $\mathcal{L}_{local}$ (Eq. 2), while the whole model is jointly optimized with the guidance of the teacher alignment module. In the second stage, we employ the decentralized DRO approach to learn a robust teacher alignment module that performs well conditioned on the worst-case local models where the feature encoder is distorted the most severely. Learning object of alleviating distortion in local triaing. After finishing the local training in each communication round, we aim to employ an extra de-centralized learning process in the client-side to optimize the alignment module on the worst-case distribution of an uncertainty set constructed by the local aligner and local data distribution, denoted as $P_{i}$ for simplicity. The uncertainty set can be written as:
$\footnotesize Q^{c}:\{D(Q||P^{c})\leq\rho\},P^{c}=\sum_{i=1}^{N}w_{i}P_{i},w\in\Delta_{N-1}$ (3)
where $N$ is the number of clients, $\Delta$ is a $(N-1)$-dim simplex, $D$ is a discrepancy measure, and $\rho$ is the uncertainty radius that measures the uncertainty degree of the unseen distributions and influences model’s generalization ability. The details of the $P_{i}$ can be seen in Appendix B. With the teacher alignment module denoted as $g^{*}$ and the aggregated encoder $f^{*}$ after the first stage training, the optimization goal can be constructed by distributions $P_{i}$ of the support set:
$\footnotesize min_{g^{*}}\ max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f^{*}),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (4)
,where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ are the losses calculated on the local clients respectively. While we learn $g^{*}$ by gradient descent, employing an alternative optimization approach, the mirror gradient descent process of optimizing w could be solved efficiently (Zhang et al., 2023b; Duchi et al., 2008):
$\footnotesize{w_{i}^{(r+1)}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (5)
@@ 4.2 Learning Unbiased Cross-modal Alignment via Distribution-based Optimization
where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$, $r$ refers to the communication turn.
Finally, we formulate $\mathcal{L}_{dro}$ to measure the performance of the teacher alignment module of alleviating distortion on the encoder. Based on Eqs 3 and 4, we update the teacher alignment module in local clients with weights w. Specifically, given a client $c$, we use $f^{*}$ learned in the first stage of the current communication turn to generate features $(\tilde{v^{c}},\tilde{l^{c}})$ which are not distorted by the local aligner during the second stage’s training. And a learnable copy of $f^{*}$ to generate distorted features $(v^{c},l^{c})$. The distorted features are input to the frozen local aligner $g$ to calculate the task loss $L_{task}$, and are also input to the teacher aligner $g^{*}$ to get calibrated representations $(z_{I}^{c*},z_{T}^{c*})$. Meanwhile, $(\tilde{v^{c}},\tilde{l^{c}})$ are input to $g^{*}$ to get representation $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$ as anchors for $(z_{I}^{c*},z_{T}^{c*})$ projected by teacher aligner using distorted features, for regularization. The learning object of the second stage could be written as:
$\footnotesize\mathcal{L}_{dro}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z_{I,i}^{c*}}-z_{I,i}^{c*})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z_{T,i}^{c*}}-z_{T,i}^{c*})^{2})+\mathcal{L}_{CL}+\mathcal{L}_{task}$ (6)
,where $\mathcal{L}_{CL}$ is the contrastive loss (Radford et al., 2021) between $\tilde{(z_{I}^{c*}},\tilde{z_{T}^{c*}})$, $\mathcal{L}_{task}$ refers to the loss function of the backbone pre-training method and $N_{c}$ refers to the number of samples in the client $c$.
@@ 5.1.1 Pre-training setup
Following (Wang et al., 2022), we utilize the MIMIC-CXR (Bigolin Lanfredi et al., 2022) dataset for biomedical vision-language pre-training. During the de-centralized pre-training process, local clients only have access to their private datasets. Following Yan et al. (2023), we employ the Latent Dirichlet Allocation (LDA) (Blei et al., 2003) to divide the MIMIC-CXR dataset based on a 5-way disease attribute to construct the heterogeneous client datasets. We set the heterogeneity degree in the LDA algorithm to be 1. The client datasets are highly heterogeneous as shown in Figure 3 and the appendix. We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 clients. Each divided dataset consists of train splits and validation splits based on the notation of the MIMIC-CXR. For the pre-training reported in the main table, we set the number of communications to 25, and we randomly sampled 50 batches of data from the client datasets for each communication. For more details, please see the appendix. For analysis experiments on out-of-distribution generalization ability, we utilize the EHRXQA (Bae et al., 2023) to evaluate the transferability of the pre-trained model to the unseen domain.
Figure 3: The descriptive statistic of the divided 5 subgroups, it can be seen that they are highly heterogeneous.
5.1.2 Downstream tasks
@@
We conduct several downstream tasks following (Wang et al., 2022) to evaluate the generalization ability of the pre-trained model. For more details, please see the appendix. Few-shot classification.We test their performance on multiple image classification benchmarks: (1) RSNA Pneumonia Detection (RSNA) (Shih et al., 2019). (2) Covidx (Wang et al., 2020). We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset, and evaluate the classification accuracy. Medical image segmentation. We conduct experiments following (Wang et al., 2022) on medical image segmentation on the RSNA (Wang et al., 2020) benchmark. We freeze the encoder and fine-tune a U-Net decoder using 1%, 10% of the training data, and then use the Dice score for evaluation. Image retrieval. To verify the pre-trained models have learned semantics and cross-modal alignments, we conduct the image retrieval task. Specifically, we utilize the validation splits of the clients for evaluation, these datasets are unseen in pre-training. We utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models.
5.1.3 Backbones and Baselines
@@
We focus on upgrading biomedical vision-language pre-training (VLP) strategies to heterogeneous federated learning scenes. Since federated heterogeneous VLP is a new topic, we adopt existing biomedical VLP baselines to federated strategies (Lu et al., 2023a). Specifically, we consider adopting simple language-image contrastive alignment (ConVIRT) (Zhang et al., 2022b; Radford et al., 2021), global-local language-image contrastive alignment(GLoRIA) (Huang et al., 2021), and Multi-Granularity Cross-modal Alignment (MGCA) (Wang et al., 2022). Since the federated multi-modal per-training is a novel setting, we adapted FedMAE (Yan et al., 2023), FedEMA (Zhuang et al., 2022), and FedX (Han et al., 2022) for comparison. These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity. For basic federated learning baselines, we consider simple averaging (FedAvg) (McMahan et al., 2017; Lu et al., 2023a), centralized training (Local), and decentralized training (Global). For fair comparisons, we combine each federated learning strategy with different biomedical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models. For fair comparisons, we re-implement the methods with the same backbones. To enable uni-modal self-supervised learning baselines to be applied to our setting, we employ the image-text contrastive loss with the same hyperparameters to our method for fairness. We utilize a ViT-base (Dosovitskiy et al., 2020) as our vision encoder and utilize the Bert-base (Devlin et al., 2018) as our text encoder. We pre-process the input following (Wang et al., 2022). We utilize an additional transformer layer from ViT-base and Bert-base as the align module for vision and language respectively.
@@ 5.2 Empirical Finding
In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our Federated distributional Robust Guidance-Based (FedRGB). We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments. We utilize the ConVIRT backbone strategy and consider the performance of the centralized pre-trained model (Global), the server model pre-trained in FedAvg(FedAvg), and the average performance of decentralized pre-trained local models (Local.avg.).
FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity. Figure 4 shows retrieval accuracies of the strategies we have considered, where the FedAvg pre-trained model achieves higher accuracies than those of decentralized ones with a large margin. However, the centralized pre-trained model acts as an upper bound and there still exists a great space for improvement for FedAvg. Table 12 in the appendix shows the detailed result. Further, we examine if updating the aggregated server model in FedAvg would result in distortion in the learned representation space. Retraining the server model on local datasets would distort its feature encoders. Figure 5 demonstrates the effect of distortion by fitting local cross-modal alignments of the corresponding client datasets. We consider a set of models that finished the 25th local iteration and their corresponding aggregated server model. While the average accuracy of local models is less than their aggregated model, the average performance of models re-trained on local clients starting from the aggregated model degrades a lot. Table 13 in the appendix shows the detailed results. Diverse information is aggregated in shallow layers that mainly extract feature semantics. We re-train the first 4 shallow layers of the aggregated server model on local clients’ datasets for 500 iteration steps. As shown in figure 6, the performance of retrained models decreases a lot, which means shallow layers that extract feature semantics are distorted by heterogeneous local datasets. Table 17 in the appendix shows the detailed results. De-centralized pre-trained deep layers that act as the align module is harmful for extracting diverse semantics. We re-train the first 4 shallow layers of the decentralized pre-trained model on the union of client datasets for 500 steps. Although the performance has successfully increased, there is still a large gap between that of the FedAvg pre-trained baselines and them as shown in Figure 7. Therefore, the frozen deep layers prevent the model from learning more diverse semantics from the scaled dataset. Table 18 in the appendix shows the detailed results. Overall, from the empirical findings, we conclude that federated multimodal pre-training is sensitive to the data heterogeneity which results in learning distorted cross-modal alignment. Simply averaging the local model weights does not solve this issue essentially. Thus, we are motivated to propose our FedRGB.
Table 2: Downstream task performance. We report the few-shot classification accuracy on Covid and RSNA, the retrieval accuracy on Chest imaGenome, and the Dice score on RSNA.
@@
Decentralized 82.27 82.83 79.45 88.25 67.44 70.75 15.51 51.12 13.62 45.97
FedX        (Han et al., 2022) 80.16 81.17 80.50 89.25 71.27 71.61 3.22 16.76 1.58 11.94
FedMAE (Yan et al., 2023) 76.97 80.66 42.25 49.00 69.48 73.38 5.05 23.91 4.34 20.86
FedEMA (Zhuang et al., 2022) 82.81 83.19 79.25 86.50 70.94 73.65 24.01 67.05 21.85 62.44
FedAvg   (Lu et al., 2023a) 83.19 83.31 78.00 88.50 69.59 71.50 28.83 72.12 25.34 66.68
FedRGB (Ours) 83.21 83.54 82.50 90.00 71.51 74.08 29.49 73.03 26.99 67.65
@@ 5.3 Main Result
Our method learns robust and enriched cross-modal alignment and has better transferability. Table 2 shows the image-text retrieval results. Here we utilize the ConVIRT as the backbone pre-training method. We report the average recall accuracy and the worst-case recall accuracy. The two sets of acc. are higher than those of FedAvg, which means our model can not only capture more cross-modal alignment but optimize the performance on the worst client dataset, which verifies the robustness and fairness of our DRO-based strategy. Table 2 shows the results of the classification and segmentation downstream tasks. In general, our method has better transferability than the traditional FedAvg strategy, most of the compared metrics of our method are higher than those of the baseline, and are closer to those of the centralized pre-trained model, which is the upper bound baseline. Other self-supervised baselines that aim to tackle the data heterogeneity have weaker performance compared to FedAvg, we analyze that this is due to the lack of the modeling of heterogeneous multi-modal alignment. Previous single-modality methods cannot be adapted directly for multi-modal data. Table 2 shows the performance of adapted self-supervised federated learning methods which focus on single-modality. From the results, we observe that our method beats these baselines by a large margin, which indicates that previous single-modality methods cannot be easily adapted directly for multi-modal data. Particularly, FedMAE and the FedX perform much worse than other federated learning strategies. We conjecture this is because both of them have ignored modeling the cross-modal alignment. Specifically, FedMAE’s masking strategies might hurt the alignment of image and text modalities, and FedX’s single-modality strategy can’t be adapted to model the image-text cross-modal interaction. Instead, we think FedAvg is still a strong baseline compared to other adapted methods, as we observed in the experiments. That’s because FedAvg doesn’t hurt the ability to learn cross-modal alignment of backbone pre-training models by adding augmented losses. Our method can be transferred to multiple multi-modal pre-training methods. Table 3 shows the downstream task performance of the MGCA and the GLoRIA backbone pre-training method when combined with our strategy. Our method has successfully upgraded MGCA and GLoRIA to adapt to the heterogeneous federated multi-modal pre-training scene, as shown in the significant improvement of the classification and segmentation downstream tasks.

### v2

@@ 4.1 Cross-Modal Alignment Binding via Guidance-Based Local Training
We combine the guidance loss with the optimization goal $\mathcal{L}_{task}$ of the initial medical vision-language pre-training method, which includes e.g. image-text contrastive loss in ConVIRT Zhang et al. (2022b). The overall loss function for guidance-based local client training can be written as:
$\scriptsize\mathcal{L}_{local}=\mathcal{L}_{guide}+\alpha\mathcal{L}_{task}$ (2)
@@
where $\alpha$ is the hyper-parameter that balances between guidance regularization and training tasks.
@@
4.2 Learning Unbiased Cross-modal Alignment via Distributionally Robust Optimization
As mentioned in section 4.1, we aim to utilize an anchor space to guide the local client training. To achieve this, we need to obtain a teacher alignment module that can capture globally unbiased and generalizable cross-domain alignment. How can we obtain such a teacher guidance model? We propose to learn such unbiased cross-modal alignment by considering the robustness across various distributions of local clients. Specifically, the teacher alignment module should encourage the model to learn well-aligned latent space while resulting in less feature distortion on the aggregated feature encoder during fitting local datasets. Furthermore, the robustness requires the teacher model to perform well on various data distributions. These insights motivate us to learn an alignment module that can well align representations and reduce feature discrepancy under various local clients, which would even well tackle the most severe distortion during local training. This motivation inspires us to model the problem as a worst-case optimization problem, which can be solved by robust optimization (DRO) Cotter et al. (2019); Namkoong & Duchi (2016). Detailed problem formulation and description can be seen in Appendix F.
Learning robust cross-modal alignment through distributionally robust optimization. Following the analysis above, we model the de-centralized pre-training framework as a two-stage DRO learning process. In the first stage, the whole model is jointly optimized with $\mathcal{L}_{local}$ (Eq. 2), and the aggregated encoder $f^{*}$ is obtained by server averaging. In the second stage, we employ the decentralized DRO to learn a robust teacher alignment module $g^{*}$ that performs well conditioned on the worst-case local models and local data distribution where the feature encoder is distorted the most severely.
We employ an extra de-centralized learning process on the client side to optimize the alignment module on the worst-case distribution of an uncertainty set Namkoong & Duchi (2016) constructed by the local clients, with the distributions denoted as $P_{i},i\in[1,...,N_{c}]$ for simplicity. The details of the uncertainty set can be seen in the Appendix Sec. F. Given teacher alignment module $g^{*}$ and the aggregated encoder $f^{*}$, the optimization goal can be constructed based on the support distribution $P_{i}$ in the uncertainty set:
$\scriptsize\min_{g^{*}}\ \max_{\textbf{w}\in\Delta_{N-1}}\sum_{i=1}^{N}w_{i}\ \mathcal{L}_{dro}(g^{*},P_{i},f^{*}),s.t.D(N\textbf{w}||\textbf{1})\leq\rho$ (3)
,where $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$ is calculated on local clients. We will illustrate it in the following part.
Optimizing teacher alignment module via alternating min-max game. To match the newly learned semantics from local datasets during local training of each communication turn $r$, we aim to continually update the teacher alignment module. For each turn $r$, we employ an alternative optimization approach. We firstly complete the local training, and optimize w via the mirror gradient descent process that could be solved efficiently Namkoong & Duchi (2016); Zhang et al. (2023b); Duchi et al. (2008):
$\scriptsize{w_{i}^{(r+1)}}=\text{Proj}_{\mathcal{Q}{c}}\left({w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}}/\sum_{i=1}^{N}{w_{i}^{(r)}e^{\gamma v_{i}^{(r)}}},\rho\right)$ (4)
@@ 4.2 Learning Unbiased Cross-modal Alignment via Distributionally Robust Optimization
where $\text{Proj}_{\mathcal{Q}{c}}$ refers to the projection to $\{w:D(N\textbf{w}||\textbf{1})\leq\rho\}$, $v_{i}$ represents the loss $\mathcal{L}_{dro}(g^{*},P_{i},f^{*})$. The weights w will be used in updating $g^{*}$ following Eq. 3 in local clients.
Finally, we propose the $\mathcal{L}_{dro}$ to learn a $g^{*}$ that can well align representations of the two modalities and reduce the discrepancy between anchor and local representations projected by $g^{*}$ and local alignment module $g$ during local training. To reduce the discrepancy, we directly minimize the distance of anchor and local representations modality-wise. Specifically, for client $c$, we use $f^{*}$ to generate features $(\tilde{I},\tilde{T})$ which are not distorted by the local aligner during the second stage’s training, and a learnable copy to generate distorted features $(I,T)$. $(\tilde{I},\tilde{T})$ and $(I,T)$ are input to $g^{*}$ to get $(z_{I}^{*},z_{T}^{*})$ and $\tilde{(z_{I}^{*}},\tilde{z_{T}^{*}})$ respectively. Then we minimize the L2 distance of them. To guarantee the alignment in latent space, we utilize multi-modal contrastive loss Radford et al. (2021); Zhang et al. (2022b) $\mathcal{L}_{CL}$ between $\tilde{(z_{I}^{*}},\tilde{z_{T}^{*}})$ to guarantee to learn well-aligned latent space. To simulate the distortion by local training, we input $(I,T)$ into $g$ to get $(z_{I},z_{T})$ to calculate $\mathcal{L}_{task}$. The learning object of the second stage of client $c$ could be written as:
$\scriptsize\mathcal{L}_{dro}=\beta(\sum_{i=1}^{N_{c}}(\tilde{z_{I,i}^{*}}-z_{I,i}^{*})^{2}+\sum_{i=1}^{N_{c}}(\tilde{z_{T,i}^{*}}-z_{T,i}^{*})^{2})+\mathcal{L}_{CL}+\mathcal{L}_{task}.$ (5)
@@ 5.1 Experiment Setting
Pre-training setup. Following Wang et al. (2022), we utilize the MIMIC-CXR Bigolin Lanfredi et al. (2022) dataset for medical vision-language pre-training. During the de-centralized pre-training process, local clients only have access to their private datasets. Following Yan et al. (2023), we employ the Latent Dirichlet Allocation (LDA) Blei et al. (2003) to divide the MIMIC-CXR dataset based on a 5-way disease attribute to construct the heterogeneous client datasets. We set the heterogeneity degree in the LDA algorithm to be 1. The client datasets are highly heterogeneous as shown in Fig. 4 and the Appendix. We divide the MIMIC-CXR into 5 heterogeneous subgroups to construct 5 clients. Each divided dataset consists of train splits and test splits based on the notation of the MIMIC-CXR. For the pre-training reported in the main table, we set the number of communications to 25, and we randomly sampled 50 batches of data from the client datasets for each communication. For more details, please see the Appendix B.1.
@@
Downstream tasks. We conduct several downstream tasks following Wang et al. (2022) to evaluate the generalization ability of the pre-trained model. For more details, please see the Appendix B.2. (1). Few-shot classification. We test their performance on multiple image classification benchmarks: (a). RSNA Pneumonia Detection (RSNA) Shih et al. (2019). (b) Covidx Wang et al. (2020). We fine-tune our pre-trained model with an additional linear layer on 1%, 10% percentage of the training dataset, and evaluate the classification accuracy. (2). Medical image segmentation. We conduct experiments following Wang et al. (2022) on medical image segmentation on the RSNA Wang et al. (2020) benchmark. We freeze the encoder and fine-tune a U-Net decoder using 1%, 10% of the training data, and then use the Dice score for evaluation. (3). Image retrieval. To verify the pre-trained models have learned semantics and cross-modal alignments, we conduct the image retrieval task. Specifically, we utilize the validation splits of the clients for evaluation, these datasets are unseen in pre-training. We utilize the top-1 recall accuracy and top-5 recall accuracy to evaluate the pre-trained models.
@@
Backbones and Baselines. We focus on enabling medical vision-language pre-training (VLP) methods to be applied to heterogeneous federated learning scenes and have considered the generalization ability of FedAID on different backbone VLP methods. Since federated heterogeneous VLP is a new topic, we adopt existing medical VLP baselines to federated strategies Lu et al. (2023a). Specifically, we consider adopting contrastive learning based methods: simple language-image contrastive alignment (ConVIRT) Zhang et al. (2022b); Radford et al. (2021), global-local language-image contrastive alignment(GLoRIA) Huang et al. (2021a), and Multi-Granularity Cross-modal Alignment (MGCA) Wang et al. (2022). Since the federated multi-modal per-training is a novel setting, we adapted FedMAE Yan et al. (2023), FedEMA Zhuang et al. (2022), FedMOON Li et al. (2021b) for comparison. These are self-supervised learning methods in the federated learning domain which also focus on tackling the data heterogeneity. For basic federated learning baselines, we consider simple averaging (FedAvg) McMahan et al. (2017); Lu et al. (2023a), decentralized training (Local), and centralized training (Global). For fair comparisons, we combine each federated learning strategy with different medical VLP methods. For baselines pre-trained in Local strategy, we report the averaged performance of the local models.
For fair comparisons, we re-implement the methods with the same backbones. To enable uni-modal self-supervised learning baselines to be applied to our setting, we employ the image-text contrastive loss with the same hyperparameters to our method for fairness. We utilize a ViT-base Dosovitskiy et al. (2020) as our vision encoder and utilize the Bert-base Devlin et al. (2018) as our text encoder. We pre-process the input following Wang et al. (2022). We utilize an additional transformer layer from ViT-base and Bert-base as the align module for vision and language respectively.
@@ 5.2 Empirical Finding
In this section, we will demonstrate our key empirical findings for federated multi-modal pre-training under heterogeneous client datasets, which actually motivates us to propose our Federated distributional Robust Guidance-Based (FedAID). We conduct experiments on the image-text retrieval task, which reflects the ability to capture vision-language alignments. Due to the page limitation, here we list our key findings, detailed results can be seen in the Appendix A:
  • FedAvg can aggregate semantics from local clients, but the performance would be damaged by data heterogeneity as shown in Fig. 5.
  • Fig. 6 shows retraining server model on local datasets would distort its feature encoders.
  • Fig. 7 shows diverse info. is aggregated in shallow layers that extract feature semantics.
  • De-centralized pre-trained deep layers that act as the align module is harmful for extracting diverse semantics as shown in Fig. 8.
Overall, from empirical findings, we conclude that federated multimodal pre-training is sensitive to data heterogeneity which results in learning distorted cross-modal alignment. Simply averaging local model weights does not solve this issue essentially. Thus, we are motivated to propose our method.
Table 1: Downstream task performance. We report the few-shot classification accuracy on Covid and RSNA, the retrieval accuracy on Chest imaGenome, and the Dice score for segmentation on RSNA.
@@
ImageNet Init ViT 79.02 80.62 63.75 83.75 60.09 67.02 3.64 16.46 2.86 14.17
FedMAE 80.81 81.06 77.50 85.75 69.48 73.38 8.29 33.03 6.92 30.49
FedEMA 82.81 83.19 79.25 86.50 70.94 73.65 24.01 67.05 21.85 62.44
FedMOON 82.51 83.21 77.75 89.25 68.99 71.28 27.85 70.88 25.34 67,19
FedAvg 83.19 83.31 78.00 88.50 69.59 71.50 28.83 72.12 25.34 66.68
FedAID (Ours) 83.21 83.54 82.50 90.00 71.51 74.08 29.49 73.03 26.99 67.65
@@ 5.3 Main Result
Our method learns robust and enriched cross-modal alignment and has better transferability. Table 1 shows the image-text retrieval results. Here we utilize the ConVIRT as the backbone pre-training method. Both average and worst-client accuracies of our method are higher than baseline’s, which means our model can not only capture more cross-modal alignment but optimize the performance on the worst client dataset, which verifies the robustness and fairness of our DRO-based strategy. Table 1 shows the results of the classification and segmentation downstream tasks. Our method beats other baseline strategies on each task and can be comparable to the centralized trained model, which acts as the upper bound. Therefore, our method has better transferability.
Previous single-modality methods cannot be adapted directly for multi-modal data. Table 1 shows the performance of adapted self-supervised federated learning methods which focus on single-modality. From the results, we observe that baselines have shown transferability on visual downstream tasks. However, in the multi-modal retrieval task, our method beats these baselines by a large margin, which indicates that previous single-modality methods cannot be easily adapted directly for multi-modal data. We conjecture this is because they have ignored modeling the cross-modal alignment. Furthermore, we think FedAvg is still a strong baseline compared to other adapted methods, as we observed in the experiments. That’s because FedAvg doesn’t hurt the ability to learn cross-modal alignment of backbone pre-training models by adding augmented losses.
Our method can be transferred to multiple multi-modal pre-training methods. Table 2 shows the downstream task performance of the MGCA and the GLoRIA backbone pre-training method when combined with our strategy. Our method has successfully upgraded MGCA and GLoRIA to adapt to the heterogeneous federated multi-modal pre-training scene, as shown in the significant improvement of the classification and segmentation downstream tasks.